a swimming pool can be filling in 11 hours if a water enters through a pipe alone, or in 21 hours if water enters through a hose alone. If water enter through both the pipe and the hose, how long will it take to fill the pool?
Answer by flame8855(424)

1/11 + 1/21= 11+21/231 = 32/231
231/32= 7.21